Calculate the dimensions of force and impulse taking velocity,density and frequency as basic quantities

Grade:11

1 Answers

Subhash Kumar Mehta
5 years ago
Let we denote velocity=v, density=d, frequency=f 
We have Force=[MLT-2]
Let 
MLT-2=vadbfc
MLT-2=[LT-1]a[ML-3]b[T-1]c
》MLT-2=MbLa-3bT-a-c
On comparing the exponents of both sides
》b=1____________(1)
》a-3b=1_________(2)
》a=4
》-a-c=-2________(3)
》c=-2
Hence dimension of force is
[F]=v4d f-2
 
Similarly you can find the dimension of Impulse.
Thanks
Regards
Subhash Kumar
St. Xavier's College, Ranchi